Status Report

NASA Space Station On-Orbit Status 14 Feb 2004

By SpaceRef Editor
February 14, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 14 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.  First weekend rest day for CDR/SO Michael Foale and FE Alexander Kaleri.

As generally on Saturdays, the crew performed the weekly 3-hr. station cleaning.  [This includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

Kaleri attended to his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.

Working off the Russian discretionary task list, Alex Kaleri performed another session with the biomedical “Pulse” experiment (MBI-9), the third for the FE.  These tests are done monthly.  [Execution of the medical cardiological assessment is controlled from the Russian payloads laptop 3, using a set respiration rate (without forced or deep breaths) and synchronizing respiration with computer-commanded “inhale” commands.  Before the experiment, arterial blood pressure is measured with the “Tensoplus” sphygmomanometer.  After the test, he reported to MCC-M/TsUP and reconfigured laptop 3 to its original settings.]

Mike Foale performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).   [When CDR Foale rebooted the Lab and Airlock (A/L) PCSs at ~5:40am, the Lab laptop did not reconnect to the C&C MDM (command & control multiplexer/demultiplexer, computer).  There are only two operational PCS laptops on ISS, neither being in the US Lab.  Mike was instructed to replace the failed hard drive (HDD) #6054 with HDD # 6029.]

The crew conducted the weekly planning conference with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

The FE performed the regular daily SOZh environmental control & life support systems maintenance in the Service Module (SM).

Foale and Kaleri conducted their regular daily physical exercise program of 2.5 hrs on TVIS treadmill, CEVIS bike, RED exerciser and VELO cycle with load trainer.

Both crewmembers had their weekly PFCs (private family conferences), Mike via S-band/audio and Sasha via VHF/audio.

The air bubble-sensitive Elektron O2 generator continues to be fickle.  It shut down again early this morning, at 3:30am EST, on failure of its primary pump, then switched to the backup pump, which failed at 3:48pm.  At 4:37am, the system was reactivated. 

The BCC (backup control center) dry run in test mode on 2/12 was only partially successful, but it served its purpose by surfacing a problem.  [The single not-for-uplink Preplanned Command Packet (PPCP) sent by Moscow’s Houston Support Group (HSG) to MCC-M/TsUP was completed, but ACT (American contingency telemetry) from 11/19/93 played back by TsUP was not successfully received in the HSG room (HSR).  The planned BCC checkout could not be performed.  Until a fix of the new CETUS1.1 software is in place, HSR will use a manual command provided by MCC-H to interface with the Russian side. The written BCC Dry Run & Checkout procedures will be modified to include this change.]

A recurring item on the Russian “job jar” task list is the continuing search, on the crew’s free time, for “missing” equipment, including some critical items required to support system operation that may be needed in the nearest future.  [Mike and Sasha are using a new list of this equipment, with pictures to facilitate search & identification.  The list calls out a BRPK condensate separation & pumping unit, two manually operated pumps, a current regulator, a hose, a BKV water treatment unit, and an IPD Draeger tubes cartridge belt.]

Preparatory to the 2/26 EVA-9, a test of the 7S/Soyuz TMA-4 MCS (motion control system), including thrusters, will be conducted on 2/16 (Monday) at 3:40am, during which the ISS will be in free drift for 40 minutes.  The station will then revert to the current LVLH YVV attitude (local vertical/local horizontal, y-axis in velocity vector) until 2/21.

As part of the new sleep shift schedule, bedtime for the crew today began at 12:30pm EST, followed by wakeup tonight at 9:00pm.  This cycle will be maintained through 2/19, to support the Soyuz ingress demo in Orlan suits on that day.  Preparatory to the EVA demo and the actual EVA-9 a week later, the crew will undergo the standard mandatory pre-EVA medical activities: on 2/16 mass and leg volume checks (MO-7, MO-8), on 2/17 urinalysis (MO-9), and on 2/18 cardiovascular evaluation during graded exercises (MO-5).

Weekly Science Update (Expedition Eight — 15th):

GASMAP:   Thanks were expressed to Mike for powering the GASMAP this week and helping to maintain the GASMAP hardware.  There will be another health check next month

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound:   Planned.

Hand Posture Analyzer (HPA:  Looking forward to future operations.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Looking forward to future operations, after the PromISS operations.
Space Acceleration Measurement System (SAMS):   Currently investigating some interesting disturbances.
Microgravity Acceleration Measurement System (MAMS):   MAMS OSS and HIRAP will capture Soyuz thruster testing on 2/16.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  PROMISS- 3 has been running flawlessly now for 12.5 days.  Protein samples are investigated in the Digital Holographic Microscope in a very stable temperature environment (25degC +/- 0.3degC).  Experimental conditions are optimal up to now.  The optical microscope also shows very stable fringe patterns.  On 2/10, a new table was uplinked to PROMISS, in order to modify the video recording rate.  This new table also contains two corrections.  There are two tape exchanges planned next week.  If necessary, payload operators may improve the science return by adjusting the stop positions for the PROMISS wheel by uploading a modified table from ground

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Renal Stone (RS):   The crew’s last in-flight data collection session will occur in April.  Thanks to the crew for their continued support with taking their daily pills at dinner time.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Pre-Increment requirements have been completed.  A second BBT (Beacon & Beacon Tester) session will be scheduled in US Lab.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Due to data dropouts during this week’s downlink (not a problem with the data transfer to the PC), another downlink will be requested.  The ground will get back to Foale with the results as soon as they are available.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Looking forward to the next set of FDI Tissue Culture runs.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new. 

Earth Science Toward Exploration Research (ESTER):  Looking forward to the upcoming sessions.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   CGBA continues to run nominally at 20C.

Miscible Fluids in Microgravity (MFMG):    Nothing new.

Educational Payload Operations (EPO):  The education demonstrations continue to be excellent.  The video will be used in a variety of ways.  The ground is looking forward to upcoming activities.

Crew Earth Observations (CEO):   A striking, detailed CEO image of Pangue Dam and reservoir on the Bio-Bio River in Chile will be published on Earth Observatory this weekend.  The development of Bio-Bio for its hydroelectric potential continues to be an intensely debated topic for Chileans.  As daylight has returned to the Northern Hemisphere and the number of target opportunities has increased, so has the volume and variety of downlinked CEO imagery from ISS.  The crew did a “fine job” acquiring the CEO target city of Addis Abeba, Ethiopia despite clouds and lack of contrasting terrain features.  Kudos for getting this challenging target!  The crew also had several close encounters with newer targets in the Libyan Desert.  A cursory survey of imagery from as recent as 2/12 suggests the crew has “nailed” a number of the ground’s coral reef targets and is also taking fascinating images of familiar features under winter conditions of snow and ice.  Camera times are good and most long lens views have good focus.

Today’s optional CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Shanghai, China (this city has just topped 20 million, which includes more than 3 million people classed as a “floating” population [mainly unregistered migrants from inland].  The station’s descending track passed just offshore: looking right on the south side of the Yangtze River estuary), Madrid, Spain (nadir pass.  The winter green-up should be well under way after the recent good winter rains.  Madrid’s population is 3 millionequal to Shanghai’s floating population), Lahore, Pakistan (nadir pass over this megacity of 6.3 million.  Lahore lies on a minor river but is easily detected at the focus point of many visible highway and railroad transport routes), and Rome, Italy (looking right towards the coast, at the south end of a line of circular lakes.  Rome’s population is 2.65 million).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 3:00am EST, 2/13).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (16A).  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (repair now completed).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 8:59am EST [= epoch]):

  • Mean altitude — 367.1 km
  • Apogee — 372.0 km
  • Perigee — 362.3 km
  • Period — 91.89 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.000719
  • Orbits per 24-hr. day — 15.67
  • Solar Beta Angle — -64.9 deg (magnitude decreasing)
  • Mean altitude loss last 24 hours — 140 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29892

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.